CH/Tech/93/12

Electronics

Predicted Spend Profile for 1994

Authors: R.A. Hunt & D. Savory Date: 10 December 1993

Introduction: Horoscope for 1994

These predictions for the period until September 1994 should be read in conjunction with the note Revised Estimates for Channel Amplifiers and Control which was issued in September 1993 (Ch/Doc/93/12).

This note will show the actual and predicted spend to complete the manufacture of the 312 amplifier control array. The figures have been arranged in the same order as the September 93 note but now under three headings A B and C. The Camac modules Pile-Up and Control are not shown in this note and now become part of a separate Data Acquisition estimate.

A shows the amount of money spent in advance where it is clear that items will definitely be required. For example standard Euro metalwork connectors and standard components.

B shows the money which could be committed in advance but where it is wise to delay the spend until after the prototype has been evaluated. At the moment 1st May 1993 is the target for evaluation of the prototype to be completed and for the design to be fixed.

C is a prediction of the money required to complete the final production. This proportion needs to be spent between May to September 1994 and should all be committed before September 1994.

Dual Amplifier 165 off A Components, card frame enclosure, connectors: £15979 B As above plus printed circuit design and production: £25736 C Production and Assembly: £9600 Walk and Accept modules 22 off

A Card frame enclosure and connectors: £477 B Components, Printed circuit layout and production, sub-circuit : Printed circuit design for sub-circuit included in prototype production. £7677 C Production and Assembly: £1860

Amplitude/Time Conversion modules 22 off A Card frame enclosure and connectors: £477 B Components, sub-circuit, printed circuit layout and production: Printed circuit design for sub-circuit included in prototype production. £9070 C Production and Assembly: £1880

Integrator 320 off A All printed circuit design costs included in prototype production. B Components, printed circuit production: £7440 C Production and Assembly: £1520

Powered Crates 12 off A Power supplies, cable assembly, enclosure metalwork, connectors: £9557 B Production and Assembly: £2400 Summary

ITEM A B C TOTAL
Dual Amplifiers £15979 £25736 £9600 £51315
Walk and Accept £477  £7677 £1860 £10014
Convertors  £477  £9070 £1880 £11427
Powered Crate  £  £9557 £2400 £11957
Sub Total  £16933 £52040 £15740 £84713

All prices exclude VAT

Separate alternative for Peak Sense or Integrate options

Sample/Hold Silena £45 320 off £14400

or

Integrator  £28 320 off £8996

All prices exclude VAT

R Hunt and D Savory 10th December 1993